OpenStudy (anonymous):

How do you change powers with negative exponents to powers with positive exponents?

OpenStudy (saifoo.khan):

Flip it.

OpenStudy (anonymous):

x^-a =1/x^a

OpenStudy (anonymous):

x^-a =1/x^a

jimthompson5910 (jim_thompson5910):

Use the idea that \[\LARGE x^{-k}=\frac{1}{x^k}\] where k is some positive number For example, say that x = 6 and k = 2, then \[\LARGE 6^{-2}=\frac{1}{6^2}\]

OpenStudy (amistre64):

put them in their place; they dont belong here ....

OpenStudy (amistre64):



OpenStudy (amistre64):

you flip them over to the other side of the fraction bar to get them postive again

OpenStudy (amistre64):

...thru the looking glass alice, thru the looking glass